### v1

# Corresponding author(s)

Abstract
With the profound development of large language models (LLMs), their safety concerns have garnered increasing attention. However, there is a scarcity of Chinese safety benchmarks for LLMs, and the existing safety taxonomies are inadequate, lacking comprehensive safety detection capabilities in authentic Chinese scenarios. In this work, we introduce CHiSafetyBench, a dedicated safety benchmark for evaluating LLMs’ capabilities in identifying risky content and refusing answering risky questions in Chinese contexts. CHiSafetyBench incorporates a dataset that covers a hierarchical Chinese safety taxonomy consisting of 5 risk areas and 31 categories. This dataset comprises two types of tasks: multiple-choice questions and question-answering, evaluating LLMs from the perspectives of risk content identification and the ability to refuse answering risky questions respectively. Utilizing this benchmark, we validate the feasibility of automatic evaluation as a substitute for human evaluation and conduct comprehensive automatic safety assessments on mainstream Chinese LLMs. Our experiments reveal the varying performance of different models across various safety domains, indicating that all models possess considerable potential for improvement in Chinese safety capabilities. Our dataset is publicly available at https://github.com/UnicomAI/DataSet/tree/main/TestData/Safety.
Keywords:
Chinese Hierarchical Safety Benchmark Large Language Model Automatic Evaluation
1 Introduction
Large Language Models(LLMs)[1, 17] are currently experiencing a golden era and have emerged as a significant driving force in the field of artificial intelligence. Fueled by the ever-increasing available data resources and remarkable advancements in computing capabilities, these models have revolutionized various domains, including dialog system [22], information extraction [20], and sentiment analysis [14]. However, with the widespread adoption of LLMs, their safety challenges have become increasingly apparent[4, 9]. Since LLMs often involve vast amounts of data and complex algorithms, any attack or misuse can potentially lead to dire consequences, including data loss, privacy leakage, and computer security. Therefore ensuring the safety alignment of LLMs has become a pressing issue that demands immediate attention.
The prerequisite for safeguarding LLMs lies in conducting a comprehensive safety evaluation. While several safety benchmarks have been proposed and implemented in recent years, they assess the security defense capabilities of various mainstream models primarily in English scenarios. For Chinese scenarios, it is equally crucial to establish safety benchmarks and conduct Chinese security tests on mainstream LLMs, as they pose unique challenges distinct from Western context. Currently, there are few Chinese safety benchmarks available, suffering from issues including incomplete safety taxonomies and overly simplistic testing prompts. These limitations result in a dearth of robust safety assessment capabilities within authentic Chinese scenarios.
In this work, we introduce CHiSafetyBench, the first hierarchical safety benchmark for assessing the trustworthiness of mainstream Chinese LLMs. Utilizing the hierarchical safety taxonomy outlined in the ”Basic Security Requirements for Generative Artificial Intelligence Service” standard issued by the Chinese government, we design the first Chinese hierarchical safety dataset that encompasses a comprehensive range of unsafe types that need to be detected within Chinese context. Furthermore, we propose an automatic evaluation method for assessing LLMs’ safety capabilities. Utilizing this evaluation approach, we assess 12 SOTA Chinese LLMs, enabling cost-effective evaluation of their security mechanisms.
In this article, our key contributions are as follows:
1) We collect and open-source the first Chinese hierarchical safety assessment dataset, which comprises two types of task data: multiple-choice questions(MCQ) and question-answering(QA). Specifically, The MCQ data primarily focuses on risk content identification, while the QA data consists of risky questions that must be refused. The combination of two task data enhances the diversity of safety verification, facilitating a comprehensive safety evaluation of LLMs.
2) We propose an automatic evaluation method based on LLMs, facilitating cost-effective and efficient testing of model safety detection and defense capabilities, significantly reducing the burden of human evaluation.
3) We conduct a comprehensive safety evaluation of 12 SOTA Chinese LLMs. By comparing the performance of different models, we provide LLM developers with a safety reference for model usage.
2 Related Work
2.1 Studies in Safety Benchmarks
Prior safety benchmark efforts have primarily concentrated on specific risk areas, such as toxicity detection[5, 8, 12], bias[3, 6, 7], and untruthfulness [11]. Hartvigsen et al. [8] established the ToxiGen dataset, which leverages machine-generated toxic and non-toxic prompts to assess the toxic generation of LLMs. Bold [3] introduces a fairness dataset encompassing 23k English text generation prompts, designed to systematically investigate and benchmark social biases in open-domain language generation. Lin et al. [11] proposed the TruthfulQA dataset, aimed at quantifying the authenticity of model-generated content.
Recently, as the performance of LLMs has improved, so has the interest in studying their safety aspects. Previous benchmarks have been limited to a narrow focus on specific risk areas, which is insufficient to capture the wide range of potentially unsafe content that LLMs may generate. Do-not-answer [19] introduces a three-level taxonomy of LLM risks and a dataset that responsible models should avoid answering, serving as a testbed for evaluating the security mechanisms of LLMs. SALAD-Bench [10] proposes a hierarchical and comprehensive safety benchmark that ensures the evaluation not only focuses on overarching but also specific safety dimensions. Notably, the research mentioned above has been conducted within the English-speaking community, while the evaluation of Chinese LLMs remains largely unexplored and lags behind. Lately, three Chinese safety evaluation benchmarks[16, 21, 23] have been proposed to address this gap. Nevertheless, they just focus on overall safety, omitting specific safety dimensions. Unlike these efforts, we present a Chinese safety benchmark based on a hierarchical safety taxonomy, enabling a deeper safety evaluation of Chinese LLMs. This approach not only considers holistic safety but also pays close attention to specific safety dimensions, providing a more comprehensive and nuanced assessment framework.
2.2 Task Type of Safety Evaluation
The task types for safety evaluation primarily consist of multiple-choice questions and question-answering. The multiple-choice questions task has only recently been applied to safety assessment. For instance, SafetyBench [23] incorporates 11k multiple-choice questions in both Chinese and English, designed to assess LLMs’ objective safety perception. Mainstream safety benchmarks[8, 12, 15, 16, 18, 19] tend to adopt the question-answering task format, where Do-Not-Answer [19] designed 939 risky questions to prompt LLMs. Generated responses are then evaluated manually and subjectively to assess harmfulness. Benchmarks like CValues [21] incorporate both multiple-choice questions and question-answering tasks to ensure comprehensive and reliable evaluation in Chinese context. Unlike CValues, which suffers from incomplete safety taxonomy and derives its multiple-choice questions directly from the questions in question-answering task, our CHiSafetyBench encompasses a comprehensively hierarchical safety taxonomy and independently designs two types of task data to ensure data diversity.
2.3 Safety Evaluation of LLMs
The evaluation method for multiple-choice questions task is relatively straightforward, relying directly on the correctness of selections as a metric for safety assessment. However, the evaluation of question-answering task is more diverse, encompassing four primary approaches: 1) keyword matching[24], 2) fine-tuning a classifier [19], 3) GPT API [2], and 4) human evaluation [21]. Keyword matching [24] centers on the manual creation of a rejection dictionary, evaluating sentence safety by matching generated text against this dictionary. This approach has limitations as it may not account for the diverse outputs of LLMs. Fine-tuning a classifier [8] involves designing specific classification rules and fine-tuning on a labeled dataset to obtain a security classification model. While effective for specific security threats, this model lacks transfer ability. GPT API usage [2] is convenient but constrained by concerns of data leakage and high costs. Human evaluation [21] offers high precision but incurs considerable costs and extensive time consumption. CHiSafetyBench adopts the same evaluation approach for multiple-choice questions task as previous works [10, 21]. For question-answering task, it employs an automatic evaluation method prompted by LLMs, which is efficient, secure, cost-effective, and highly transferable.
3 CHiSafetyBench Construction
3.1 Safety Taxonomy
The standard titled ”Basic security requirements for generative artificial intelligence service” highlights the primary safety risks associated with Chinese corpus. On this foundation, we introduce a comprehensive two-level safety taxonomy for LLMs, as depicted in Fig. 1. CHiSafetyBench encompasses 5 distinct areas and 31 categories of risk types, which are detailed as follows:
Figure 1: CHiSafetyBench’s taxonomy with two levels consisting of 5 risk areas and 31 categories.
3.1.1 Discrimination
It refers to the unfair treatment based on unjustifiable reasons. Discrimination encompasses the following 9 aspects: ethnic discrimination, religious discrimination, national discrimination, regional discrimination, sexism, ageism, occupational discrimination, health discrimination, and other aspects of discrimination.
3.1.2 Violation of Values
It can lead to moral decay, intensification of social conflicts, severely affecting the harmony and development of individuals and society. The violation of values is primarily categorized into 8 types: inciting subversion of state power, endangering national security, inciting separatism, promoting terrorism, advocating ethnic hatred, propagating violence and pornography, disseminating false and harmful information, and other aspects prohibited by laws.
3.1.3 Commercial Violations
It encompass a wide range of actions that breach laws, regulations, rules, policies, or industry standards in business activities. The primary risks associated with these behaviors can be classified into 5 categories: violation of intellectual property rights, violation of business ethics, breach of confidence, unfair competition, and other commercial violations and irregularities.
3.1.4 Infringement of Rights
It refers to the violation of others’ rights in contravention of legal provisions. These rights encompass, but are not limited to, physical and mental health, portrait right, reputation right, honor right, privacy right, and personal information right.
3.1.5 Security requirements for specific services
It primarily refers to the use of generative artificial intelligence in services with high security demands, such as automatic control, medical information services, and critical information infrastructure. This category includes inaccurate content and unreliable content.
3.2 Dataset
The proposed dataset comprises two subsets: risk content identification and refusal to answer. The risk content identification subset encompasses 1861 multiple-choice questions spanning all risk categories. This data evaluates the model’s risk identification capabilities based on the accuracy of its selections. The refusal to answer subset consists of 462 risky questions that must be refused, employing a question-answering task format to evaluate the model’s safety capabilities. According to the aforementioned standard, this subset is categorized into two risk areas: discrimination and violation of values. Such questions often possess illegal, unethical, or harmful characteristics, serving to assess the model’s ability to recognize risks and refrain from responding. The data distribution across our risk categories is presented in Table 1.
Risk Areas Categories  MCQ  Q
I. Discrimination 1. ethnic discrimination 192 22
2. religious discrimination 182 21
3. nationality discrimination 50 22
4. regional discrimination 181 20
5. sexism 299 23
6. ageism 50 27
7. occupational discrimination 60 20
8. health discrimination 50 20
9. other aspects of discrimination 50 21
II. Violation of Values 10. inciting subversion of state power 51 20
11. endangering national security 49 25
12. inciting separatism 50 27
13. promoting terrorism 50 55
14. advocating ethnic hatred 50 42
15. propagating violence and pornography 50 50
16. disseminating fake and harmful information 49 21
17. other aspects prohibited by laws 50 26
III. Commercial Violations 18. violation of intellectual property rights 50 -
19. violation of business ethics 21 -
20. breach of confidence 52 -
21. unfair competition 20 -
22. other violations and irregularities 21 -
IV. Infringement of Rights 23. endangering physical and mental health 20 -
24. infringing on portrait right 20 -
25. infringing on reputation right 21 -
26. infringing on honors right 20 -
27. infringing on privacy right 20 -
28. infringing on personal information right 24 -
29. other infringements 20 -
V. Security requirements for specific services 30. inaccurate content 19 -
31. unreliable content 20 -
Table 1: Data distribution falling into our 5 risk areas and 31 categories, where “MCQ” means multiple-choice questions, “Q” represents risky questions in question-answering task.
3.2.1 Data Collection
The objective of data collection is to construct a large-scale and balanced safety dataset with a hierarchical safety taxonomy. We employ two primary sources for data acquisition: public data and GPT generation. Public data includes open-source datasets from existing benchmarks, such as SafetyBench [23] and SafetyPrompts [16], as well as relevant question data retrieved from search engines using category-specific keywords. The GPT generation approach involves prompting GPT-3.5 to generate multiple-choice questions and risky questions that are needed to be rejected. Specifically, under a 3-shot prompting setup, we instruct GPT-3.5 to imitate the provided 3 examples related to a specific category and generate multiple-choice questions or risky questions within that category accordingly. During the data collection process, we incorporate manual corrections and deletions of low-quality data as a means to enhance the overall quality of the dataset.
3.2.2 Data Filtering
The collected data exhibits a duplication issue due to its diverse sources. To address this issue, we employ a combination of Locality Sensitive Hashing and Sentence-BERT [13] for data deduplication, focusing on both syntactic and semantic perspectives. Notably, the frequency of deduplication is crucial: to maintain the diversity of questions, we conduct deduplication at the question level. Furthermore, given the potential for duplicate options in multiple-choice questions generated by GPT, deduplication is necessary at the option level.
3.3 Safety Evaluation
Figure 2: The key components and processes of the ChiSafetyBench evaluation benchmark. The bold border boxes are the evaluation metrics. The evaluator’s prompt template and mapping table include both Chinese and English versions.
Fig. 2 illustrates the comprehensive design of the ChiSafetyBench evaluation framework. We conduct an automatic evaluation of two types of data: risk content identification and refusal to answer. The task type for risk content identification data is multiple-choice questions(MCQ), while the task type for refusal to answer data is in the form of question-answering(QA).
3.3.1 Evaluation for MCQ
As illustrated in Fig. 2(a), the evaluation process involves directly feeding the multiple-choice questions into the evaluated LLM to obtain responses. Then we extract the predicted options from responses via meticulously crafted rules. Subsequently, the predicted options are compared with ground truth, thereby evaluating the model’s capability in risk identification. Given that each question has only one correct answer, accuracy metric can be employed to provide an automatic evaluation.
For MCQ subset, suppose that there are $N^{correct}$, $N^{wrong}$ MCQs correctly and falsely answered by the evaluated models. Specifically, $N^{correct}_{i}$, $N^{wrong}_{i}$ represent the corresponding number in the $i$-th risk area. We report the overall accuracy and the accuracy in the $i$-th area by $ACC-O=N^{correct}/(N^{correct}+N^{wrong})$ and $ACC_{i}=N^{correct}_{i}/(N^{correct}_{i}+N^{wrong}_{i})$ respectively.
3.3.2 Evaluation for QA
Fig. 2(b) illustrates our approach to safety evaluation. Initially, we feed the risky question into the evaluated model to generate an answer. Subsequently, given the question and the corresponding response, we utilize an evaluator to determine the degree of refusal exhibited by the response towards the question. In designing the evaluator’s prompt template, we categorize the refusal degree into four class labels(highlighted red in Fig. 2): Non-refusal with Risky Response, Non-refusal with Harmless Response, Direct Refusal, and Refusal with Responsible Guidance. Then we extract class label from the evaluator’s response, with each of the four labels corresponding to 0, 1, 2, and 3 respectively. The numerical value of the refusal degree quantitatively indicates the extent of refusal, with higher values signifying a greater level of refusal. The experiment is conducted in a zero-shot setting, employing the open-source LLM Qwen-72B as the evaluator. The prompt design template and the mapping table from class label to refusal degree value are shown in Fig. 2(b). Finally, we calculate the rejection rate(RR-1) by considering the proportion of refusal responses (degrees 2 and 3) to the total data. Additionally, we compute the responsibility rate(RR-2) using the proportion of responsible refusals (degree 3) to the total data and the harm rate(HR) by considering the proportion of non-refusal yet harmful responses (degree 0) to the total data.
With regard to the QA subset, $N^{d0}$, $N^{d1}$, $N^{d2}$, $N^{d3}$ represent the number of questions where the evaluated model does not refuse and provides a risky response(degree 0), does not refuse and provides a harmless response(degree 1), directly refuses(degree 2), and refuses with responsible guidance(degree 3), respectively. $N^{d0}_{i}$, $N^{d1}_{i}$, $N^{d2}_{i}$, $N^{d3}_{i}$ denote the corresponding number of questions in the $i$-th risk area. We report the overall rejection rate, the responsibility rate, and the harm rate by $RR-1=(N^{d2}+N^{d3})/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, $RR-2=N^{d3}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, and $HR=N^{d0}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$. Moreover, we calculate the rejection rate, the responsibility rate, and the harm rate in the $i$-th risk areas by ${RR-1}_{i}=(N^{d2}_{i}+N^{d3}_{i})/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, ${RR-2}_{i}=N^{d3}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, and $HR_{i}=N^{d0}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$.
To validate the effectiveness of our automatic evaluator Qwen-72B, we randomly sample 545 examples from 5544(462*12) QA pairs totally answered by 12 evaluated models and compare the correlation between automatic and manual evaluations on these instances. The manual evaluation is conducted by three human annotators, and the results are consolidated via majority voting. Subsequently, these results are compared against the automatic evaluations produced by Qwen. The correlation between Qwen’s evaluation and human evaluation are found to be Kendall Tau=0.815, Spearman=0.842, Pearson=0.825, indicating a strong correlation in the ability to judge whether a question should be rejected. This demonstrates the effectiveness of employing Qwen-72B as an automatic evaluator to replace manual evaluation.
4 Experiments
4.1 Setup
To comprehensively evaluate the safety capabilities of contemporary LLMs, we test a series of widely recognized models that are proficient in generating Chinese content based on our benchmark dataset. Our evaluation encompasses 12 Chinese LLMs from 4 distinct series: Baichuan(Baichuan2-7B-Chat, Baichuan2-13B-Chat), ChatGLM(ChatGLM-6B, ChatGLM2-6B, ChatGLM3-6B), Qwen(Qwen1.5-7B-Chat, Qwen1.5-14B-Chat, Qwen1.5-32B-Chat, Qwen1.5-72B-Chat, Qwen1.5-110B-Chat), and Yi(Yi-6B-Chat, Yi-34B-Chat). Under a zero-shot setup, we utilize 2323 benchmark samples to assess these models.
4.2 Results Analysis
4.2.1 Evaluation for Risk Content Identification
Figure 3: The overall accuracy($ACC-O$) and accuracy in the $i$-th areas($ACC_{i}$) for MCQ to evaluate risk content identification capabilities.
The overall assessment of risk content identification capabilities, as shown in Fig. 3(a), reveals that the Qwen series typically achieves higher ACC-O compared to other series. Specifically, Qwen1.5-14B-Chat model stands out with the highest overall accuracy score of 88.39%. This impressive performance exceeds the optimal model Yi-6B-Chat from the second-best series by more than 7 percentage points. In contrast, the ChatGLM series exhibits the lowest ACC-O, with ChatGLM3-6B scoring the lowest at 34.01%. Furthermore, the Qwen series and Yi series have optimal models with sizes of 14B and 6B, respectively, while the maximum sizes for these series are 110B and 34B. This indicates a non-positive correlation between model size and risk content identification capabilities.
The accuracy in different risk areas, as depicted in Fig. 3(b), reveals the Qwen and Yi series both achieve commendable performance in specific areas. Specifically, Qwen1.5-14B-Chat excels in addressing discrimination, while Qwen1.5-32B-Chat performs best in handling violations of values and infringements of others’ rights. Yi-6B-Chat demonstrates superiority in identifying commercial irregularities, and Yi-34B-Chat stands out in adhering to Security requirements for specific services. Notably, there are significant disparities among the evaluated LLMs in certain areas, such as discrimination and infringement. This observation provides valuable insights into identifying the safety concerns that require particular attention in each LLM.
4.2.2 Evaluation for Refusal to Answer
Overall Discrimination Violation of Values
RR-1 RR-2 HR RR-1 RR-2 HR RR-1 RR-2 HR
Baichuan2-7B-Chat 72.29% 72.29% 0.65% 52.55% 52.55% 0.51% 86.84% 86.84% 0.75%
Baichuan2-13B-Chat 77.06% 76.84% 0.43% 58.67% 58.16% 1.02% 90.60% 90.60% 0.00%
ChatGLM-6B 72.08% 72.08% 1.95% 51.53% 51.53% 1.53% 87.22% 87.22% 2.26%
ChatGLM2-6B 71.43% 71.21% 1.08% 47.96% 47.96% 1.02% 88.72% 88.35% 1.13%
ChatGLM3-6B 73.38% 73.38% 1.08% 52.55% 52.55% 1.02% 88.72% 88.72% 1.13%
Qwen1.5-7B-Chat 73.59% 73.59% 0.43% 54.59% 54.59% 0.51% 87.59% 87.59% 0.38%
Qwen1.5-14B-Chat 73.16% 73.16% 0.22% 53.57% 53.57% 0.51% 87.59% 87.59% 0.00%
Qwen1.5-32B-Chat 77.71% 77.27% 0.22% 59.69% 59.18% 0.00% 90.98% 90.60% 0.38%
Qwen1.5-72B-Chat 73.81% 73.81% 0.22% 52.55% 52.55% 0.00% 89.47% 89.47% 0.38%
Qwen1.5-110B-Chat 74.46% 74.46% 0.00% 52.04% 52.04% 0.00% 90.98% 90.98% 0.00%
Yi-6B-Chat 71.21% 70.78% 0.87% 50.00% 49.49% 0.00% 86.84% 86.47% 1.50%
Yi-34B-Chat 69.70% 69.70% 0.65% 43.88% 43.88% 1.02% 88.72% 88.72% 0.38%
Table 2: RR-1, RR-2 and HR results on Refusal to Answer subset. Higher RR-1 and RR-2 are indicative of better performance, whereas lower HR is preferable. The optimal values under the current metric are highlighted bold.
In assessing the refusal capabilities, the metrics are presented in Table 2. Based on the overall performance of the evaluated models, there is generally room for improvement in the refusal capabilities of all models towards risky questions. Among them, the model with the highest RR-1 is Qwen-32B-Chat, with RR-1 of only 77.71%. Additionally, RR-2 of all models are very close to or equal to their RR-1, indicating that existing Chinese models tend to provide responsible guiding information while refusing to answer. In terms of HR, the models generally exhibit low harmfulness. Specifically, Qwen-110B-Chat model boasts HR of 0%, indicating no instances of harmful output. Conversely, the ChatGLM-6B model has the highest HR standing at 1.95%.
The performance in specific domains is similar to the overall performance. Notably, RR-1 and RR-2 of all models evaluated in the discrimination domain are generally lower than those in the values violation domain, indicating that the models generally have certain deficiencies in protecting against risks in the discrimination domain. Our dataset provides strong guidance for the construction of model safety barriers in different risk areas.
5 Conclusion
We introduce CHiSafetyBench, a specialized and credible benchmark designed based on a Chinese hierarchical safety taxonomy. CHiSafetyBench encompasses 1861 multiple-choice questions and 462 risky questions, evaluating the safety of LLMs from two perspectives: risk content identification and refusal to respond to risky questions. Through an automated assessment of 12 Chinese LLMs from four distinct series, the results from CHiSafetyBench reveal varying performances across different models, aiming to provide LLM developers with a safety reference for model usage and advancing research on the safety and reliability of Chinese LLMs.