### v2

# CHiSafetyBench: A Chinese Hierarchical Safety Benchmark for Large Language Models

Abstract
With the profound development of large language models(LLMs), their safety concerns have garnered increasing attention. However, there is a scarcity of Chinese safety benchmarks for LLMs, and the existing safety taxonomies are inadequate, lacking comprehensive safety detection capabilities in authentic Chinese scenarios. In this work, we introduce CHiSafetyBench, a dedicated safety benchmark for evaluating LLMs’ capabilities in identifying risky content and refusing answering risky questions in Chinese contexts. CHiSafetyBench incorporates a dataset that covers a hierarchical Chinese safety taxonomy consisting of 5 risk areas and 31 categories. This dataset comprises two types of tasks: multiple-choice questions and question-answering, evaluating LLMs from the perspectives of risk content identification and the ability to refuse answering risky questions respectively. Utilizing this benchmark, we validate the feasibility of automatic evaluation as a substitute for human evaluation and conduct comprehensive automatic safety assessments on mainstream Chinese LLMs. Our experiments reveal the varying performance of different models across various safety domains, indicating that all models possess considerable potential for improvement in Chinese safety capabilities. Our dataset is open-sourced on https://github.com/UnicomAI/UnicomBenchmark/tree/main/
CHiSafetyBench.
1 Introduction
Large Language Models(LLMs)(Achiam et al. 2023; Touvron et al. 2023) are currently experiencing a golden era and have emerged as a significant driving force in the field of artificial intelligence. Fueled by the ever-increasing available data resources and remarkable advancements in computing capabilities, these models have revolutionized various domains, including dialog system (Zhang et al. 2023a), information extraction (Xu et al. 2023a), and sentiment analysis (Rusnachenko, Golubev, and Loukachevitch 2024). However, with the widespread adoption of LLMs, their safety challenges have become increasingly apparent(Dong et al. 2024; Huang et al. 2023). Since LLMs often involve vast amounts of data and complex algorithms, any attack or misuse can potentially lead to dire consequences, including data loss, privacy leakage, and computer security. Therefore ensuring the safety alignment of LLMs has become a pressing issue that demands immediate attention.
Figure 1: CHiSafetyBench’s taxonomy with two levels consisting of 5 risk areas and 31 categories.
The prerequisite for safeguarding LLMs lies in conducting a comprehensive safety evaluation. While several safety benchmarks have been proposed and implemented in recent years, they assess the security defense capabilities of various mainstream models primarily in English scenarios. For Chinese scenarios, it is equally crucial to establish safety benchmarks and conduct Chinese security tests on mainstream LLMs, as they pose unique challenges distinct from Western context. Currently, there are few Chinese safety benchmarks available, suffering from issues including incomplete safety taxonomies and overly simplistic testing prompts. These limitations result in a dearth of robust safety assessment capabilities within authentic Chinese scenarios.
In this work, we introduce CHiSafetyBench, the first Chinese Hierarchical Safety Benchmark for assessing the trustworthiness of mainstream Chinese LLMs. Utilizing the hierarchical safety taxonomy outlined in the ”Basic Security Requirements for Generative Artificial Intelligence Service” standard issued by the Chinese government, we design a Chinese hierarchical safety dataset that encompasses a comprehensive range of unsafe types that need to be detected within Chinese context. Furthermore, we propose an automatic evaluation method for assessing LLMs’ safety capabilities. Utilizing this evaluation approach, we assess 10 SOTA Chinese LLMs, enabling cost-effective evaluation of their security mechanisms. In this article, our key contributions are as follows:
  • We introduce CHiSafetyBench, a hierarchical safety evaluation benchmark tailored for Chinese safety scenarios. It encompasses a two-tiered structure, comprising 5 major areas and 31 categories. Additionally, we offer multi-level fine-grained evaluations tailored to this hierarchical structure, ensuring comprehensive assessment.
  • We create and open-source the first Chinese hierarchical risk assessment dataset, which comprises two types of task data: multiple-choice questions(MCQ) and question-answering(QA). The MCQ data primarily focuses on identifying risk content, while the QA data incorporates risky questions that must be refused. The integration of these two types of task data enhances the diversity of safety verification.
  • We innovatively introduce risk questions with conversational history into the QA task, in addition to single-turn risk questions. The inclusion of dialogue history better aligns with realistic interactions and poses more complex challenges for testing the security boundaries of LLMs.
  • We propose an LLM-based automatic evaluation method that rivals human assessment, facilitating cost-effective and efficient testing of model safety detection and defense capabilities, significantly reducing the burden of human evaluation.
  • We conduct a comprehensive safety evaluation of 10 SOTA Chinese LLMs. By comparing the performance of different models, we provide LLM developers with a safety reference for model usage.
Risk Areas Categories  MCQ  Q
I. Discrimination 1. ethnic discrimination 50 26
2. religious discrimination 50 27
3. nationality discrimination 50 29
4. regional discrimination 50 24
5. sexism 50 31
6. ageism 50 33
7. occupational discrimination 60 26
8. health discrimination 50 25
9. other aspects of discrimination 50 24
II. Violation of Values 10. inciting subversion of state power 51 27
11. endangering national security 49 31
12. inciting separatism 50 35
13. promoting terrorism 50 65
14. advocating ethnic hatred 50 47
15. propagating violence and pornography 50 55
16. disseminating fake and harmful information 49 26
17. other aspects prohibited by laws 50 32
III. Commercial Violations 18. violation of intellectual property rights 50 -
19. violation of business ethics 51 -
20. breach of confidence 52 -
21. unfair competition 50 -
22. other violations and irregularities 51 -
IV. Infringement of Rights 23. endangering physical and mental health 50 -
24. infringing on portrait right 50 -
25. infringing on reputation right 51 -
26. infringing on honors right 50 -
27. infringing on privacy right 50 -
28. infringing on personal information right 54 -
29. other infringements 50 -
V. Security requirements for specific services 30. inaccurate content 49 -
31. unreliable content 50 -
Table 1: Data distribution falling into our 5 risk areas and 31 categories, where “MCQ” means multiple-choice questions, “Q” represents risky questions in question-answering task. In the question-answering task, the data ratio between single-turn questions without dialogue history and questions with dialogue history is approximately four to one.
2 Related Work
Studies in Safety Benchmarks
Prior safety benchmark efforts primarily concentrate on specific risk areas, such as toxicity detection(Gehman et al. 2020; Hartvigsen et al. 2022; Lin et al. 2023), bias(Dhamala et al. 2021; Han et al. 2022; Han, Baldwin, and Cohn 2023), and untruthfulness (Lin, Hilton, and Evans 2022). Hartvigsen et al. (Hartvigsen et al. 2022) establishes the ToxiGen dataset, which leverages machine-generated toxic and non-toxic prompts to assess the toxic generation of LLMs. Bold (Dhamala et al. 2021) introduces a fairness dataset encompassing 23k English text generation prompts, designed to systematically investigate and benchmark social biases in open-domain language generation. Lin et al. (Lin, Hilton, and Evans 2022) proposed the TruthfulQA dataset, aimed at quantifying the authenticity of model-generated content. HarmfulQA(Bhardwaj and Poria 2023) utilizes an adversarial dialogue framework involving red and blue teams to generate 7.3K English harmful conversations.
Previous benchmarks have been limited to a narrow focus on specific risk areas, which is insufficient to capture the wide range of potentially unsafe content that LLMs may generate. Therefore, Do-Not-Answer (Wang et al. 2024) introduces a three-level taxonomy of LLM risks and a dataset that responsible models should avoid answering, serving as a testbed for evaluating the security mechanisms of LLMs. SALAD-Bench (Li et al. 2024) proposes a hierarchical and comprehensive safety benchmark that ensures the evaluation not only focuses on overarching but also specific safety dimensions. Notably, existing hierarchical safety benchmarks predominantly focus on English scenarios, whereas our proposed hierarchical benchmark is centered on Chinese scenarios, aiming to provide a more in-depth assessment of Chinese LLM’s safety.
Concurrently, existing Chinese benchmarks, such as SafetyPrompts (Sun et al. 2023), SafetyBench(Zhang et al. 2023b), CValues(Xu et al. 2023b) and SC-Safety(Xu et al. 2023c), lack a fine-grained, hierarchical approach to safety evaluation, often concentrating solely on overall or a few broad categories of safety. In contrast, our benchmark provides a hierarchical structure that attends not only to overall safety but also delves into specific safety dimensions.
Task Type of Safety Evaluation
The task types for safety evaluation primarily consist of multiple-choice questions and question-answering. The multiple-choice questions task has only recently been applied to safety assessment. For instance, SafetyBench (Zhang et al. 2023b) incorporates 11k multiple-choice questions in both Chinese and English, designed to assess LLMs’ objective safety perception. Mainstream safety benchmarks(Hartvigsen et al. 2022; Lin et al. 2023; Shen et al. 2023; Sun et al. 2023; Wang et al. 2023, 2024) tend to adopt the question-answering task format, where Do-Not-Answer (Wang et al. 2024) designed 939 risky questions to prompt LLMs. Generated responses are then evaluated manually and subjectively to assess harmfulness. SC-Safety(Xu et al. 2023c) is a multi-round adversarial benchmark that comprises 4912 open-ended questions. In contrast to SafetyBench, which only incorporates multiple-choice questions, Do-Not-Answer with merely risky questions, and SC-Safety featuring non-open-source multi-round questions, CHiSafetyBench encompasses both MCQ and QA task types, with all data being openly accessible. Moreover, its question-answering tasks encompass both single-round questions and those with conversational histories, providing a comprehensive assessment of model performance.
Benchmarks like CValues (Xu et al. 2023b) incorporate both multiple-choice questions and question-answering tasks to ensure comprehensive and reliable evaluation in Chinese context. Unlike CValues, which suffers from incomplete safety taxonomy and derives its multiple-choice questions directly from the questions in question-answering task, our CHiSafetyBench encompasses a comprehensively hierarchical safety taxonomy and independently designs two types of task data to ensure data diversity.
Safety Evaluation of LLMs
The evaluation method for multiple-choice questions task is straightforward, relying directly on the correctness of selections as a metric for safety assessment. However, the evaluation of question-answering task is more diverse, encompassing four primary approaches: 1) keyword matching(Zou et al. 2023), 2) fine-tuning a classifier (Wang et al. 2024), 3) GPT API (Deng et al. 2023), and 4) human evaluation (Xu et al. 2023b). Keyword matching (Zou et al. 2023) centers on the manual creation of a rejection dictionary, evaluating sentence safety by matching generated text against this dictionary. This approach has limitations as it may not account for the diverse outputs of LLMs. Fine-tuning a classifier (Hartvigsen et al. 2022) involves designing specific classification rules and fine-tuning on a labeled dataset to obtain a security classification model. While effective for specific security threats, this model lacks transfer ability. GPT API usage (Deng et al. 2023) is convenient but constrained by concerns of data leakage and high costs. Human evaluation (Xu et al. 2023b) offers high precision but incurs considerable costs and extensive time consumption. CHiSafetyBench adopts the same evaluation approach for multiple-choice questions task as previous works (Li et al. 2024; Xu et al. 2023b). For question-answering task, it employs an automatic evaluation method prompted by LLMs, which is efficient, secure, cost-effective, and highly transferable.
Figure 2: Examples of multiple choice questions and question-answering subset. Please note that the English translations are only for illustrative purposes, while the original data only contains Chinese.
3 CHiSafetyBench Construction
Safety Taxonomy
The standard titled ”Basic security requirements for generative artificial intelligence service” highlights the primary safety risks associated with Chinese corpus. On this foundation, we introduce a comprehensive two-level safety taxonomy for LLMs, as depicted in Fig. 1. CHiSafetyBench encompasses 5 distinct areas and 31 categories of risk types, which are detailed as follows:
Discrimination
It refers to the unfair treatment based on unjustifiable reasons. Discrimination encompasses the following 9 aspects: ethnic discrimination, religious discrimination, national discrimination, regional discrimination, sexism, ageism, occupational discrimination, health discrimination, and other aspects of discrimination.
Violation of Values
It can lead to moral decay, intensification of social conflicts, severely affecting the harmony and development of individuals and society. The violation of values is primarily categorized into 8 types: inciting subversion of state power, endangering national security, inciting separatism, promoting terrorism, advocating ethnic hatred, propagating violence and pornography, disseminating false and harmful information, and other aspects prohibited by laws.
Commercial Violations
It encompass a wide range of actions that breach laws, regulations, rules, policies, or industry standards in business activities. The primary risks associated with these behaviors can be classified into 5 categories: violation of intellectual property rights, violation of business ethics, breach of confidence, unfair competition, and other commercial violations and irregularities.
Infringement of Rights
It refers to the violation of others’ rights in contravention of legal provisions. These rights encompass, but are not limited to, physical and mental health, portrait right, reputation right, honor right, privacy right, and personal information right.
Security requirements for specific services
It primarily refers to the use of generative artificial intelligence in services with high security demands, such as automatic control, medical information services, and critical information infrastructure. This category includes inaccurate content and unreliable content.
Dataset
The proposed dataset comprises two subsets: risk content identification and refusal to answer. The risk content identification subset encompasses 1567 multiple-choice questions spanning all risk categories. This data assesses the model’s risk identification capabilities based on the accuracy of its selections. The refusal to answer subset contains 563 risky questions that must be declined, utilizing a question-answering task format to evaluate the model’s safety capabilities. Notably, the refusal to answer subset not only includes single-turn questions without conversational history but also incorporates questions with multi-turn dialogue history. The inclusion of multi-turn dialogue data better evaluates the model’s proficiency in handling complex safety issues. According to the aforementioned standard, the refusal to answer subset is categorized into two risk areas: discrimination and violation of values. Such questions often possess illegal, unethical, or harmful characteristics, serving to assess the model’s ability to recognize risks and refrain from responding. The data distribution across our risk categories is presented in Table 1.
Data Collection
Figure 3: The key components and processes of the ChiSafetyBench evaluation benchmark. The bold border boxes are the evaluation metrics. The evaluator’s prompt template and mapping table include both Chinese and English versions.
The objective of data collection is to construct a large-scale and balanced safety dataset with a hierarchical safety taxonomy. Given that LLMs inherently generate security data that they can handle due to their inherent security mechanisms, our research avoids automatic generation methods for large models and instead adopts a dual approach that utilizes open resources and manual writing to obtain data.
In constructing the MCQ task subset, we initially use the open-source dataset SafetyBench (Zhang et al. 2023b) as the baseline, aligning it with our predefined safety taxonomy. For categories that are insufficiently covered or data-scarce, we supplement the question bank by searching search engines with category-related keywords. The category keywords are listed in Appendix A.
For single-turn questions without dialogue history, we employ manual writing to ensure the effective exclusion of low-quality data. As for questions with multi-turn dialogue history, we take harmfulQA  (Bhardwaj and Poria 2023) as the baseline, translating and meticulously aligning it within our safety taxonomy. For categories with uneven data distribution, we also adopt a manual writing strategy to provide targeted supplementation, thereby achieving a comprehensive and balanced dataset.
Data Filtering
The collected data exhibits a duplication issue due to its diverse sources. To address this issue, we employ a combination of Locality Sensitive Hashing and Sentence-BERT (Reimers and Gurevych 2019) for data deduplication, focusing on both syntactic and semantic perspectives and conduct deduplication at the question level. After data filtering, the final subsets of MCQ and QA were obtained, with data examples presented in Fig. 3. The entire version of the examples is provided in Appendix B.
Safety Evaluation
Fig. 3 illustrates the comprehensive design of the ChiSafetyBench evaluation framework. We conduct an automatic evaluation of two types of data: risk content identification and refusal to answer. The task type for risk content identification data is multiple-choice questions(MCQ), while the task type for refusal to answer data is in the form of question-answering(QA).
Evaluation for MCQ
As illustrated in Fig. 3(a), the evaluation process involves directly feeding the multiple-choice questions into the evaluated LLM to obtain responses. Then we extract the predicted options from responses via meticulously crafted rules. Subsequently, the predicted options are compared with ground truth, thereby evaluating the model’s capability in risk identification. Given that each question has only one correct answer, accuracy metric can be employed to provide an automatic evaluation.
For MCQ subset, suppose that there are $N^{correct}$, $N^{wrong}$ MCQs correctly and falsely answered by the evaluated models. Specifically, $N^{correct}_{i}$, $N^{wrong}_{i}$ represent the corresponding number in the $i$-th risk area. We report the overall accuracy and the accuracy in the $i$-th area by $ACC-O=N^{correct}/(N^{correct}+N^{wrong})$ and $ACC_{i}=N^{correct}_{i}/(N^{correct}_{i}+N^{wrong}_{i})$ respectively.
Evaluation for QA
Figure 4: The overall accuracy($ACC-O$) and accuracy in the $i$-th areas($ACC_{i}$) for MCQ to evaluate risk content identification capabilities.
Fig. 3(b) illustrates our approach to safety evaluation. Initially, we feed the risky question into the evaluated model to generate an answer. Subsequently, given the question and the corresponding response, we utilize an evaluator to determine the degree of refusal exhibited by the response towards the question. In designing the evaluator’s prompt template, we categorize the refusal degree into four class labels(highlighted red in Fig. 3): Non-refusal with Risky Response, Non-refusal with Harmless Response, Direct Refusal, and Refusal with Responsible Guidance. Then we extract class label from the evaluator’s response, with each of the four labels corresponding to 0, 1, 2, and 3 respectively. The numerical value of the refusal degree quantitatively indicates the extent of refusal, with higher values signifying a greater level of refusal. The experiment is conducted in a zero-shot setting, employing the open-source LLM Qwen-72B as the evaluator. The prompt design template and the mapping table from class label to refusal degree value are shown in Fig. 3(b). Finally, we calculate the rejection rate(RR-1) by considering the proportion of refusal responses (degrees 2 and 3) to the total data. Additionally, we compute the responsibility rate(RR-2) using the proportion of responsible refusals (degree 3) to the total data and the harmfulness rate(HR) by considering the proportion of non-refusal yet harmful responses (degree 0) to the total data.
With regard to the QA subset, $N^{d0}$, $N^{d1}$, $N^{d2}$, $N^{d3}$ represent the number of questions where the evaluated model does not refuse and provides a risky response(degree 0), does not refuse and provides a harmless response(degree 1), directly refuses(degree 2), and refuses with responsible guidance(degree 3), respectively. $N^{d0}_{i}$, $N^{d1}_{i}$, $N^{d2}_{i}$, $N^{d3}_{i}$ denote the corresponding number of questions in the $i$-th risk area. We report the overall rejection rate, the responsibility rate, and the harmfulness rate by $RR-1=(N^{d2}+N^{d3})/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, $RR-2=N^{d3}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, and $HR=N^{d0}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$. Moreover, we calculate the rejection rate, the responsibility rate, and the harmfulness rate in the $i$-th risk areas by ${RR-1}_{i}=(N^{d2}_{i}+N^{d3}_{i})/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, ${RR-2}_{i}=N^{d3}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, and $HR_{i}=N^{d0}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$.
To validate the effectiveness of our automatic evaluator Qwen-72B, we randomly sample 563 examples from 5630(563*10) QA pairs totally answered by 10 evaluated models and compare the correlation between automatic and manual evaluations on these instances. The manual evaluation is conducted by three human annotators, and the results are consolidated via majority voting. Subsequently, these results are compared against the automatic evaluations produced by Qwen. The correlation between Qwen’s evaluation and human evaluation are found to be Kendall Tau=0.815, Spearman=0.842, Pearson=0.825, indicating a strong correlation in the ability to judge whether a question should be rejected. This demonstrates the effectiveness of employing Qwen-72B as an automatic evaluator to replace manual evaluation.
Overall Discrimination Violation of Values
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$
Baichuan2-7B-Chat 64.30% 64.30% 3.02% 49.80% 49.80% 1.63% 75.47% 75.47% 4.09%
Baichuan2-13B-Chat 68.03% 67.85% 3.37% 52.65% 52.24% 2.45% 79.87% 79.87% 4.09%
ChatGLM3-6B 65.36% 65.19% 3.91% 49.80% 49.39% 2.04% 77.36% 77.36% 5.35%
Qwen1.5-7B-Chat 67.32% 67.32% 1.07% 53.06% 53.06% 0.82% 78.30% 78.30% 1.26%
Qwen1.5-14B-Chat 66.43% 66.25% 2.13% 51.84% 51.43% 0.82% 77.67% 77.67% 3.14%
Qwen1.5-32B-Chat 69.63% 69.27% 1.60% 56.33% 55.92% 0.41% 79.87% 79.56% 2.52%
Qwen1.5-72B-Chat 68.03% 68.03% 1.78% 53.47% 53.47% 0.82% 79.25% 79.25% 2.52%
Qwen1.5-110B-Chat 69.98% 69.98% 0.36% 53.47% 53.47% 0.00% 82.70% 82.70% 0.63%
Yi-6B-Chat 62.70% 62.17% 4.62% 45.71% 44.90% 3.27% 75.79% 75.47% 5.66%
Yi-34B-Chat 63.06% 63.06% 3.73% 43.27% 43.27% 2.86% 78.30% 78.30% 4.40%
Table 2: RR-1, RR-2 and HR results on the refusal to answer subset. The optimal values under the current metric are highlighted bold.
4 Experiments
Experimental Setup
To comprehensively evaluate the safety capabilities of contemporary LLMs, we test a series of widely recognized models that are proficient in generating Chinese content based on our benchmark dataset. Our evaluation encompasses 10 Chinese LLMs from 4 distinct series: Baichuan(Baichuan2-7B-Chat, Baichuan2-13B-Chat), ChatGLM(ChatGLM3-6B), Qwen(Qwen1.5-7B-Chat, Qwen1.5-14B-Chat, Qwen1.5-32B-Chat, Qwen1.5-72B-Chat, Qwen1.5-110B-Chat), and Yi(Yi-6B-Chat, Yi-34B-Chat). Under a zero-shot setup, we utilize 2030 benchmark samples to assess these models.
Evaluation for Risk Content Identification
The overall assessment of risk content identification capabilities, as shown in Fig. 8(a), reveals that the Qwen series typically achieves higher ACC-O compared to other series. Specifically, Qwen1.5-72B-Chat stands out with the highest overall accuracy score of 91.13%. This impressive performance exceeds the optimal model Yi-6B-Chat from the second-best series by more than 5 percentage points. In contrast, the ChatGLM series exhibits the lowest ACC-O, with ChatGLM3-6B scoring the lowest at 41.16%. Furthermore, the Qwen series and Yi series have optimal models with sizes of 72B and 6B, respectively, while the maximum sizes for these series are 110B and 34B. This indicates a non-positive correlation between model size and risk content identification capabilities.
The accuracy in different risk areas, as depicted in Fig. 8(b), reveals the Qwen and Yi series both achieve commendable performance in specific areas. Specifically, Qwen1.5-14B-Chat excels in addressing discrimination, while Qwen1.5-32B-Chat performs best in handling violation of values, commercial irregularities and security requirements for specific services. Qwen1.5-72B-Chat demonstrates superiority in identifying infringements of others’ rights. Notably, there are significant disparities among the evaluated LLMs in certain areas, such as discrimination and infringement. This observation provides valuable insights into identifying the safety concerns that require particular attention in each LLM. The specific ACC of 10 models across 31 fine-grained categories can be found in Appendix C.1.
Evaluation for Refusal to Answer
As indicated in Table 2, the Qwen series consistently exhibit stronger abilities to refuse risky questions compared to other series, whereas the Yi series demonstrates notable deficiencies in managing refusal risks.
In terms of the overall performance of the models, Qwen1.5-110B-Chat and Yi-6B-Chat respectively demonstrate the best and worst capabilities in managing refusal risks across the three evaluation metrics. Specifically, regarding the refusal capabilities, even the best-performing Qwen-110B-Chat achieves a RR-1 of merely 69.98%, underscoring the prevalent room for improvement in all models’ ability to mitigate refusal risks. Additionally, the RR-2 of all models are very close or equal to their RR-1, indicating that existing Chinese LLMs tend to provide responsible guiding information while declining to answer. As for HR, these models generally exhibit low levels of potential harm, with Qwen-110B-Chat boasting a HR as low as 0.36%, whereas Yi-6B-Chat has the highest HR at 4.62%.
The performance in specific domains is similar to the overall performance. Notably, RR-1 and RR-2 of all models evaluated in the discrimination domain are generally lower than those in the values violation domain, indicating that the models generally have certain deficiencies in protecting against risks in the discrimination domain. Our dataset provides valuable guidance for the construction of model safety barriers in different risk areas. Detailed performances across 17 fine-grained categories are presented in Appendix C.2.
Ablation Study
To affirm the indispensability of multi-turn dialogue data within the refusal to answer subset, we further divide this subset into two subsets: risky questions without and with conversational history. Subsequently, we evaluate the overall RR-1, RR-2, and HR metrics for these two subsets separately, and the results are summarized in Table 3. The detailed performance of 10 models in response to risk questions without and with conversational history can be found in Appendix C.3.
Table 3 reveals that all models encounter significant obstacles in rejecting risky questions with conversational history. Compared to single-round questions, the model’s RR-1 for risk questions with dialogue history decreased by 24.96% to 50.33%, highlighting the notable deficiencies of current Chinese LLMs in handling risky dialogues. When evaluating the LLMs’ responsible guidance capabilities, the peak RR-2 score for single-turn QA reaches 77.27%, demonstrating their effectiveness in guiding under single-turn risky situations. However, in multi-turn dialogue environments, the peak RR-2 score plummets to 49.50%, indicating a significant degradation of the models’ guidance abilities in complex dialog scenarios. Furthermore, the HR for single-turn questions are generally low, with the highest being only 1.98%, whereas they sharply increase to a maximum of 26.73% in multi-turn dialogues, further corroborating the limitations of the models in managing multi-turn risky dialogues. In summary, existing models exhibit notable shortcomings in handling complex risky dialogues, confirming the necessity of incorporating multi-turn data.
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ Baichuan2-7B-Chat 72.29%/27.72% 72.29%/27.72% 0.65%/13.86% Baichuan2-13B-Chat 77.06%/26.73% 76.84%/26.73% 0.43%/26.73% ChatGLM3-6B 73.38%/28.71% 73.38%/27.72% 1.08%/16.83% Qwen1.5-7B-Chat 73.59%/38.61% 73.59%/38.61% 0.43%/3.96% Qwen1.5-14B-Chat 73.16%/35.64% 73.16%/34.65% 0.22%/10.89% Qwen1.5-32B-Chat 77.71%/32.67% 77.27%/32.67% 0.22%/7.92% Qwen1.5-72B-Chat 73.81%/41.58% 73.81%/41.58% 0.22%/8.91% Qwen1.5-110B-Chat 74.46%/49.50% 74.46%/49.50% 0.00%/1.98% Yi-6B-Chat 71.21%/23.76% 70.78%/22.77% 0.87%/21.78% Yi-34B-Chat 69.70%/32.67% 69.70%/32.67% 0.65%/17.82%
Table 3: The performance between risky questions without and with conversational history on the refusal to answer subset. From left to right, the values reported are as follows: risky questions without and with conversational history.
5 Conclusion
We introduce CHiSafetyBench, a specialized and credible benchmark designed based on a Chinese hierarchical safety taxonomy. CHiSafetyBench encompasses 1567 multiple-choice questions and 563 risky questions, evaluating the safety of LLMs from two perspectives: risk content identification and refusal to respond to risky questions. Through an automated assessment of 10 Chinese LLMs from four distinct series, the results from CHiSafetyBench reveal varying performances across different models, aiming to provide LLM developers with a safety reference for model usage and advancing research on the safety and reliability of Chinese LLMs.
Appendix A Appendix A. Search Keywords for Multiple Choice Questions
Figure 5: Keyword examples for 31 category, with ’-’ indicating that the MCQ data for this category is sourced from SafetyBench and does not require keywords to search engines. English version of these keywords are also provided.
Appendix B Appendix B. Examples
Figure 6: Entire examples of multiple choice questions and question-answering subset. Please note that the English translations are only for illustrative purposes, while the original data only contains Chinese.
Appendix C Appendix C. Experiments Details
C.1 ACC for MCQ across 31 Risk Categories
Figure 7: ACC for 10 models across 31 risk categories on the MCQ subset.
C.2 Detail Performance on Refusal of Answer Subset
Figure 8: RR-1 for 10 models across 17 risk categories on the refusal to answer subset.
Figure 9: RR-2 for 10 models across 17 risk categories on the refusal to answer subset.
Figure 10: HR for 10 models across 17 risk categories on the refusal to answer subset.
C.3 Evaluation Metrics of Risky Questions without and with Conversational History
Overall Discrimination Violation of Values
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$
Baichuan2-7B-Chat 72.29% 72.29% 0.65% 52.55% 52.55% 0.51% 86.84% 86.84% 0.75%
Baichuan2-13B-Chat 77.06% 76.84% 0.43% 58.67% 58.16% 1.02% 90.60% 90.60% 0.00%
ChatGLM3-6B 73.38% 73.38% 1.08% 52.55% 52.55% 1.02% 88.72% 88.72% 1.13%
Qwen1.5-7B-Chat 73.59% 73.59% 0.43% 54.59% 54.59% 0.51% 87.59% 87.59% 0.38%
Qwen1.5-14B-Chat 73.16% 73.16% 0.22% 53.57% 53.57% 0.51% 87.59% 87.59% 0.00%
Qwen1.5-32B-Chat 77.71% 77.27% 0.22% 59.69% 59.18% 0.00% 90.98% 90.60% 0.38%
Qwen1.5-72B-Chat 73.81% 73.81% 0.22% 52.55% 52.55% 0.00% 89.47% 89.47% 0.38%
Qwen1.5-110B-Chat 74.46% 74.46% 0.00% 52.04% 52.04% 0.00% 90.98% 90.98% 0.00%
Yi-6B-Chat 71.21% 70.78% 0.87% 50.00% 49.49% 0.00% 86.84% 86.47% 1.50%
Yi-34B-Chat 69.70% 69.70% 0.65% 43.88% 43.88% 1.02% 88.72% 88.72% 0.38%
Table 4: RR-1, RR-2 and HR results of risky questions without conversational history on the refusal to answer subset.
Overall Discrimination Violation of Values
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$
Baichuan2-7B-Chat 27.72% 27.72% 13.86% 38.78% 38.78% 6.12% 17.31% 17.31% 21.15%
Baichuan2-13B-Chat 26.73% 26.73% 16.83% 28.57% 28.57% 8.16% 25.00% 25.00% 25.00%
ChatGLM3-6B 28.71% 27.72% 16.83% 38.78% 36.73% 6.12% 19.23% 19.23% 26.92%
Qwen1.5-7B-Chat 38.61% 38.61% 3.96% 46.94% 46.94% 2.04% 30.77% 30.77% 5.77%
Qwen1.5-14B-Chat 35.64% 34.65% 10.89% 44.90% 42.86% 2.04% 26.92% 26.92% 19.23%
Qwen1.5-32B-Chat 32.67% 32.67% 7.92% 42.86% 42.86% 2.04% 23.08% 23.08% 13.46%
Qwen1.5-72B-Chat 41.58% 41.58% 8.91% 57.14% 57.14% 4.08% 26.92% 26.92% 13.46%
Qwen1.5-110B-Chat 49.50% 49.50% 1.98% 59.18% 59.18% 0.00% 40.38% 40.38% 3.85%
Yi-6B-Chat 23.76% 22.77% 21.78% 28.57% 26.53% 16.33% 19.23% 19.23% 26.92%
Yi-34B-Chat 32.67% 32.67% 17.82% 40.82% 40.82% 10.20% 25.00% 25.00% 25.00%
Table 5: RR-1, RR-2 and HR results of risky questions with conversational history on the refusal to answer subset.